Circular No. 6057

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994W IN NGC 4041 R. J. Cumming and P. Lundqvist, Stockholm Observatory; and W.

P. S. Meikle, Imperial College, London, report: “A spectrogram (range 428-800 nm, resolution 0.12 nm) of SN 1994W was taken with the 4.2m William Herschel Telescope (+ ISIS) at La Palma by M. Breare and M. Azzaro on Aug. 13.9 UT. The Hα profile shows the same features as described on IAUC 6046 and is unchanged since a high-resolution (0.03-nm) Isaac Newton Telescope spectrogram was taken by E. J. Zuiderwijk on July 31. In addition to the features reported on IAUC 6046, we identify broad (FWHM 2000 km/s) emission in He I at 447.1 and 706.5 nm, and narrow P-Cyg lines with absorption minima at ∼ −500 km/s (relative to the velocity of H II emission from the galaxy near the supernova) in Mg II (448.1 nm), He I (587.6 and 667.8 nm), Si II (634.7 and 637.1 nm), and O I (777.3 nm). An additional broad (FWHM 1500 km/s) emission feature at 792 nm may be due to a blend of Mg II (789.6 and 787.7 nm). The constancy of the H I profiles and the presence of narrow P-Cyg absorption suggest that the supernova is exciting a massive, dense circumstellar shell, rather than a radiatively accelerated stellar wind. We expect the fastest supernova ejecta to reach this shell soon, causing the dramatic disappearance of the narrow absorptions as in SN 1984E (cf. IAUC 6048), broadening of the Balmer lines, and perhaps x-ray and radio emission. Continued observations at all wavelengths are encouraged.”

UZ BOOTIS Further visual magnitude estimates (cf. IAUC 6056): Aug. 16.84 UT,

[14.3 (L. Szentaskó, Veresegyház, Hungary); 17.490, 12.9 (Y. Takenaka, Kyoto, Japan); 17.849, 12.0 (J. Pietz, Erftstadts, Germany); 18.188, 11.6 (R. Royer, Lakewood, CA); 18.477, 11.7 (T. Kato, Ouda Station, Japan; CCD + V filter); 18.819, 12.1 (Szentaskó); 18.869, 12.0 (P. Schmeer, Bischmisheim, Germany).

COMET MACHHOLZ (1994o) Total visual magnitude estimates: Aug. 14.42 UT, 9.4 (D. Machholz,

Colfax, CA, 0.25-m reflector); 16.080, 9.7 (K. Hornoch, Lelekovice, Czech Republic, 0.13-m reflector); 16.23, 10.5 (C. E. Spratt, Victoria, BC, 0.20-m reflector); 17.23, 10.3 (Spratt).

1994 August 19 Daniel W. E. Green